jigglypuff314 (jigglypuff314):

Evaluate the indefinite integral \[\int\limits\frac{ 2y+3 }{ (y+7)^{3} }dy\] I used u substitution u = y+7 and got \[\frac{ -2(y-4) }{ (y+7)^{2} } + C\] would that be correct? :3

jigglypuff314 (jigglypuff314):

My work... Evaluate the indefinite integral \[\int\limits\frac{ 2y+3 }{ (y+7)^{3} }dy\] u = y+7 -> y = u-7 du = 1 ∫(2(u-7)+3)(u^-3)du ∫(2u-11)(u^-3)du ∫(2u^-2 - 11u^-3) du -2u^-1 + 22u^-2 + C -2u^-2 (u-11) + C (-2 / (y+7)^2) (y+7-11) + C \[\frac{ -2(y-4) }{ (y+7)^{2} } + C\]

hartnn (hartnn):

what have you done after \((2(u-7)+ 3 ) u^{-3}\) ?

jigglypuff314 (jigglypuff314):

distribute the 2 add like terms distribute the u^-3 then take the integral ?

hartnn (hartnn):

well, you didn't do it correctly

hartnn (hartnn):

2(u-7)+3 = 2u -14 +3

hartnn (hartnn):

isn't it ?

jigglypuff314 (jigglypuff314):

yeah so 2u - 11

jigglypuff314 (jigglypuff314):

then distribute the u^-3 ?

hartnn (hartnn):

oh, yeah that step is correct.

OpenStudy (anonymous):

\[\int\limits \frac{ 2y+3 }{ \left( y+7 \right)^3 }dy=\int\limits \frac{ 2y+14-11 }{\left( y+7 \right)^3 }dy\] \[=2\int\limits \left( y+7 \right)^{-2}dy-11\int\limits \left( y-7 \right)^{-3}dy\] =?

jigglypuff314 (jigglypuff314):

ummm... so what'd I do wrong?

hartnn (hartnn):

i don't think you went wrong anywhere, rechecking the resubstitution part

hartnn (hartnn):

integral of 11u^(-3) is ?

hartnn (hartnn):

11 u^(-2)/ (-2) right ??

hartnn (hartnn):

finally found the error after going it over 2-3 times :P

jigglypuff314 (jigglypuff314):

ohhhh okay, thanks :)

hartnn (hartnn):

did you get it too ? whats your new answer ?

jigglypuff314 (jigglypuff314):

\[\frac{ -2 }{ y+7 } + \frac{ 11 }{ 2(y+7)^{2} } + C\]?

jigglypuff314 (jigglypuff314):

-11(-1/2) u^-2 to 11 u^-2 / 2 ?

hartnn (hartnn):

-4 (y+7) + 11 = -4y -28 +11 = -4y -17 you're correct :)

hartnn (hartnn):

\(\frac{ -2 }{ y+7 } + \frac{ 11 }{ 2(y+7)^{2} } + C \quad \huge \checkmark\)

jigglypuff314 (jigglypuff314):

okay, THANK YOU!!! :D <3

hartnn (hartnn):

welcome ^_^

OpenStudy (anonymous):

correction i wrote y-7,it is y+7

jigglypuff314 (jigglypuff314):

yeah, got that ^_^ thanks for your help!

OpenStudy (anonymous):

yw